Hi, I am having trouble working out this question, help would be much appreciated.
An electrically charged particle generates an electric field around it. The strength, E, of the electric field (in NC^-1) is directly proportional to the size of the charge of the particle, Q (in Coulombs, C), and inversely proportional to the square of the distance from it, d in metres.
(i). Set up the proportionality relationship
This is what I think the formula is but not sure if it is correct?
E (NC^-1)= k * (Q/D^2)
And then theirs question (ii)
(ii). Find the constant of proportionality if the electric field strength 3.00 cm from a 3.5 uC charge is 3.5 x 10^7 NC^-1
This is what I've worked out so far not sure if it is correct though?
3.5 * 10^7 = K * Q/(0.03)^2
3.5 * 10^7 = K * Q/(0.09)

Answer by ikleyn(52787)

.
Hi, I am having trouble working out this question, help would be much appreciated.
An electrically charged particle generates an electric field around it. The strength, E, of the electric field (in NC^-1)
is directly proportional to the size of the charge of the particle, Q (in Coulombs, C),
and inversely proportional to the square of the distance from it, d in metres.
(i). Set up the proportionality relationship
This is what I think the formula is but not sure if it is correct?
E (NC^-1)= k * (Q/D^2)
And then theirs question (ii)
(ii). Find the constant of proportionality if the electric field strength 3.00 cm from a 3.5 uC charge is 3.5 x 10^7 NC^-1
This is what I've worked out so far not sure if it is correct though?
3.5 * 10^7 = K * Q/(0.03)^2
3.5 * 10^7 = K * Q/(0.09)
--------------------------------------------------

Hi, I am having trouble working out this question, help would be much appreciated. 

An electrically charged particle generates an electric field around it. The strength, E, of the electric field (in NC^-1) 
is directly proportional to the size of the charge of the particle, Q (in Coulombs, C), 
and inversely proportional to the square of the distance from it, d in metres. 

(i). Set up the proportionality relationship 
This is what I think the formula is but not sure if it is correct? 

E (NC^-1)= k * (Q/D^2)     <------ it is correct, but replace D by d

And then theirs question (ii) 

(ii). Find the constant of proportionality if the electric field strength 3.00 cm from a 3.5 uC charge is 3.5 x 10^7 NC^-1 

This is what I've worked out so far not sure if it is correct though? 

3.5 * 10^7 = K * Q/(0.03)^2
3.5 * 10^7 = K * Q/(0.09)       <----- (0.03)^2 = 0.0009